Ralph Corporation is a job-order costing company that uses activity-based costing to apply overhead to jobs. The following overhead activities were budgeted for the year: Activity Cost Driver Amount of Driver Setups $240,000 number of setups 6,000 Purchasing $160,000 number of parts 20,000 Other overhead $300,000 direct labor hours 100,000 The following information about the jobs was given for the month of April. Job 01 Job 02 Job 03 Job 04 Balance on April 1 $64,900 $40,800 $30,500 Direct materials $54,000 $37,900 $25,000 $11,000 Direct labor $80,000 $38,500 $43,000 $21,000 Number of setups 40 10 30 200 Number of parts 300 80 400 500 Direct labor hours 5,000 2,400 5,200 1,200 By April 30, Jobs 02 and 03 were completed and sold. The remaining jobs were still in process. Question: What is the cost of goods manufactured?
Answer #1

Manufacturing overhead assignment rate on activity based costing basis:

Set-ups rate = Set-up cost / number of setups = $240000/6000 = $40 per setup

Parts Purchase rate = Purchasing cost / Number of parts = $160000 / 20000 = $8 per part

Other Overhead rate = overhead / total direct labor hours = $300000 / 100000 DLH = $3 per DLH

JOB 1 JOB 2 JOB 3 JOB 4 Total
Opening Balance 64900 40800 30500 -
Direct Material 54000 37900 25000 11000
Direct Labor 80000 38500 43000 21000
Manufacturing Overhead:
Set-up costs 40*40=1600 10*40=400 30*40=1200 200*40=8000
Part purchase costs 300*8=2400 80*8=640 400*8=3200 500*8=4000
Other Overheads cost 5000*3=15000 2400*3=7200 5200*3=15600 1200*3=3600
Total costs 217900 125440 118500 47600
Jobs Incomplete or WIP -217900       -47600
COST OF GOODS MANUFACTURED 0 125440 118500 0 $243940